OpenStudy (anonymous):

is this true sin60 csc60=1

OpenStudy (thomas5267):

By definition, \(\csc(x)=\dfrac{1}{\sin(x)}\).

OpenStudy (imstuck):

\[\sin(60)=\frac{ \sqrt{3} }{ 2 }\]The csc would then be opposite that (flip the fraction) and then multiply. The terms will then all cancel each other out to come out to 1, so yes it does equal 1.

OpenStudy (anonymous):

thank you so much

OpenStudy (imstuck):

You're very welcome!